OpenStudy (kj4uts):

How do you solve 3^(-4+3x)=27 and 5^(2x-4)=392? Round to two decimal places. Thank you!

jhonyy9 (jhonyy9):

like a first step rewrite 27 in an exponential form of 3

OpenStudy (kj4uts):

so is that 27^3?

OpenStudy (anonymous):

\[ 3^1=3\\3^2=9\\3^3=27 \]Therefore \(3^{-4+3x} = 3^3\implies -4+3x=3\).

OpenStudy (kj4uts):

Oh I see so do I now add -4 + 4 to the 3 on the other side?

OpenStudy (anonymous):

Yes

OpenStudy (kj4uts):

3x=7 and 7 / 3 = 2.33

OpenStudy (kj4uts):

I am going to try the second one but how do you determine the exponential easily because the second one is 392

OpenStudy (anonymous):

You have to use this property: \[ b =a^{\log_a(b)} \]So: \[ 392 = 5^{\log_5(392)} \]

OpenStudy (anonymous):

\[ 5^{2x-4}=5^{\log_5(392)} \implies 2x-4 =\log_5(392) \]

OpenStudy (anonymous):

You will have to use a calculator for this.

OpenStudy (kj4uts):

log base 5 (392) = 3.71

OpenStudy (anonymous):

Then solve for \(x\).

OpenStudy (kj4uts):

oh I see plus 4 = 7.71 / 2 = 3.85

OpenStudy (kj4uts):

thank you